# From Discrete to Continuous Binary Best-Response Dynamics: Discrete Fluctuations Almost Surely Vanish with Population Size

### v2

@@ VII-G Proof of Theorem 2
As for Part 1, the instability of coordinator-driven equilibria is shown by providing an initial condition from which for a small $\epsilon$, there exists no $\delta$ to satisfy the definition of stability. The stability of clean-cut and anticoordinator-driven equilibria is first shown for the abstract dynamics using Theorem 1 in [50]. Then, the definition of asymptotic stability is used to show that of clean-cut and anticoordinator-driven equilibria for the semicontinuous population dynamics. Part 2 of the theorem is shown by using the definition of limit set.
@@
Part 1) Regarding the instability analysis of a coordinator-driven equilibrium point, say $\mathbf{q}^{*}_{k-1,k}=\mathbf{o}^{ij}$, consider the following initial condition $\bm{x}(0)=\mathbf{o}^{ij}+\delta^{*}\mathbf{e}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}$ for an arbitrarily small $\delta^{*}\in(0,\min\{\tau^{\prime}_{j+1},\tau_{i}\}-\tau^{\prime}_{j})$. The corresponding abstract state $x(0)$ will then belong to $(\tau^{\prime}_{j},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$. Pick $\epsilon$ to be equal to $0.5(\min\{\tau^{\prime}_{j+1},\tau_{i},{\eta}_{i}+{\eta}_{j}^{\prime}\}-\tau^{\prime}_{j})$. In view of (5), at such $x(0)$, $\bm{\mathcal{V}}(\bm{x})$ is a singleton. Moreover, considering the structure of $\mathbf{o}^{ij}$, at $x(0)$ we have $\dot{x}_{l}=0$ for $l\in\{1,\ldots,\mathsf{p}+\mathsf{p}^{\prime}-j,\mathsf{p}+\mathsf{p}^{\prime}-j+2,\ldots,\mathsf{p}+\mathsf{p}^{\prime}\}$. Indeed, the equation $\dot{x}_{l}=0$ holds true as long as $x\in(\tau^{\prime}_{j},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$. On the other hand, $x(0)>\tau^{\prime}_{j}$ and in view of (12), we have $\dot{x}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)=\rho^{\prime}_{j}-x_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)>0$, for all $t\leq t^{\prime}$, where $t^{\prime}=\inf\{t\geq 0|x_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)=\min\{\alpha_{i,j},{\rho}^{\prime}_{j}\}\}$, where $\alpha_{i,j}=\min\{\tau^{\prime}_{j+1},\tau_{i}\}-({\eta}_{i}+{\eta}_{j-1})$. If $\min\{\alpha_{i,j},{\rho}^{\prime}_{j}\}=\alpha_{i,j}$, then $x(t^{\prime})=\min\{\tau^{\prime}_{j+1},\tau_{i}\}$. Otherwise, $x(t^{\prime})={\eta}_{j}+{\eta}_{i}$. In either case, we observe that $|x(t^{\prime})-\tau_{j}^{\prime}|$ is greater than the selected value for $\epsilon$, i.e., $0.5(\min\{\tau^{\prime}_{j+1},\tau_{i},{\eta}_{i}+{\eta}_{j}^{\prime}\}-\tau^{\prime}_{j})$, and the value of $\delta^{*}$ does not impact the value of $x(t^{\prime})$. This indicates instability of ${o}^{ij}$ and in turn $\mathbf{o}^{ij}$. The stability properties of the other two types of equilibrium points, are first shown for the abstract dynamics. Let $V_{k}(x):\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}=\frac{1}{2}(x-q^{*}_{k})^{2}$, where $q^{*}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}$ if $q^{*}_{k}$ is clean-cut, i.e., $q^{*}_{k}=c^{ij}$ and $q^{*}_{k}=\tau_{i}$ if it is anticoordinator-driven, i.e., $q^{*}_{k}=a^{ij}$. The set-valued derivative of $V_{k}(x)$ w.r.t (9) will be $\mathcal{D}(x)=\{(x-q^{*}_{k})\nu|\nu\in\mathcal{X}(x)\}$ [50]. We also have $\bar{f}(x)=\max\{y|y\in\mathcal{X}(x)\}<0$ for $x\in(q^{*}_{k},q^{*}_{k,k+1})$ and $\underline{f}(x)=\min\{y|y\in\mathcal{X}(x)\}>0$ for $x\in(q^{*}_{k-1,k},q^{*}_{k})$ (refer to the proof of 1). Hence, $\max\mathcal{D}(x)<0$ for $x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})/\{q^{*}_{k}\}$. According to [50, Theorem 1], $q^{*}_{k}$ is a strongly asymptotically stable equilibrium point for (9). Also the basin of attraction of $q^{*}_{k}$, will be $(q^{*}_{k-1,k},q^{*}_{k,k+1})$. As for $q^{*}_{1}$ (resp. $q^{*}_{\mathtt{Q}}$), we have $\bar{f}(0)>0$ (resp. $\bar{f}(1)<0$) and this case can be handled similarly which results in $\mathcal{A}(q^{*}_{1})=[0,q^{*}_{1,2})$ (resp. $\mathcal{A}(q^{*}_{\mathtt{Q}})=(q^{*}_{\mathtt{Q}-1,\mathtt{Q}},1]$). Now, we show the asymptotic stability of $\mathbf{q}^{*}_{k}$ when it is an anticoordinator-driven equilibrium point, i.e., $\mathbf{q}^{*}_{k}=\mathbf{a}^{ij}$. The case where $\mathbf{q}^{*}_{k}$ is clean-cut can be handled similarly. We show that for every $\varepsilon>0$, there exists a $\delta>0$ such that if $|\bm{x}(0)-\mathbf{a}^{ij}|<\delta$, then $|\bm{x}(t)-\mathbf{a}^{ij}|<\varepsilon$ for all $t\geq 0$, and $|\bm{x}(t)-\mathbf{a}^{ij}|\rightarrow 0$ as $t\rightarrow\infty$. Consider a small enough $\epsilon>0$, such that at an $\epsilon$-neighborhood of $a^{ij}$, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, $\mathcal{V}_{l}(\bm{x})$ is a singleton and $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ or for $l$ satisfying both $\mathsf{p}<l$ and $0<\mathsf{p}+\mathsf{p}^{\prime}+1-l\leq j$. For the remaining $l$’s, $\dot{x}_{l}=-x_{l}$. From the stability of the abstract dynamics, the existence of $\delta_{x}>0$ follows such that $|x(0)-a^{ij}|<\delta_{x}$ results in $|x(t)-a^{ij}|<\epsilon$ for $t\geq 0$. As a result, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, the final value of $x_{l}$ is $a^{ij}_{l}$ and consequently $x_{l}$ will approach $a^{ij}_{l}$ exponentially, i.e.,
@@
On the other hand, starting from $\delta_{x}$-neighborhood of $a^{ij}$, in view of (17), we have $|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<\big{(}(\mathsf{p}+\mathsf{p}^{\prime})\max_{l}|x_{l}(0)-a^{ij}_{l}|+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|e^{-t}\big{)}^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|^{2}e^{-2t}$ and, in turn, $|\bm{x}(t)-\mathbf{a}^{ij}|$ $<\max_{l}|x_{l}(0)-a^{ij}_{l}|\big{(}\mathsf{p}+\mathsf{p}^{\prime}+(\mathsf{p}+\mathsf{p}^{\prime}-1)e^{-t})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)e^{-2t}$$\big{)}^{0.5}.$ If $\delta$ is less than $\delta_{x}/(\mathsf{p}+\mathsf{p}^{\prime})$, then in view of the relation between Euclidean norm and Norm-1, $\sum_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}|x_{l}(0)-a_{l}^{ij}|$ will be less than $\delta_{x}/\sqrt{\mathsf{p}+\mathsf{p}^{\prime}}$. and, in turn, $|x(0)-a^{ij}|<\delta_{x}$. Hence, $\mathbf{a}^{ij}$ is stable because by choosing $\delta=\min\{\frac{\delta_{x}}{\mathsf{p}+\mathsf{p}^{\prime}},\frac{\varepsilon}{\sqrt{(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}+(\mathsf{p}+\mathsf{p}^{\prime}-1))^{2}+(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}-1)}\}},$ we have $|x(t)-a^{ij}|<\epsilon$ for all $t\geq 0$, and consequently $|\bm{x}(t)-\mathbf{a}^{ij}|<\varepsilon$. As for asymptotic stability, we have $|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<(|x(t)-a^{ij}|+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|e^{-t})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|^{2}e^{-2t}$, for $t\geq 0$. Since each term in the right hand side approaches zero, we have $\lim|\bm{x}(t)-\mathbf{a}^{ij}|\rightarrow 0$ as $t\rightarrow\infty$. Now, we obtain the basin of attraction of $\mathbf{q}^{*}_{k}$. (i) It is straightforward to show that if $x(0)$ falls in the basin of attraction $q^{*}_{k}$, the abstract dynamics will converge to an arbitrary small neighborhood of $q^{*}_{k}$ in a finite time. Now, depending on the types of $q^{*}_{k}$, two cases may happen: Case 1) $q^{*}_{k}=c^{ij}$, i.e., $q^{*}_{k}$ is clean-cut, then let $t=t_{0}$ be the time instant at which the abstract state $x$ enters a small enough $\epsilon$-neighborhood of $q^{*}_{k}$, such that for $l\leq i$ or $l\geq\max\{\mathsf{p}+\mathsf{p}^{\prime}-j+1,\mathsf{p}+1\}$, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining $l$’s, we have $\dot{x}_{l}=-x_{l}$. Then for $t>t_{0}$, the asymptotic value of $x_{l}$ for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$ is the same as the $l$-th entry of $\mathbf{q}^{*}_{k}$. Moreover, each $x_{l}$ approaches $c^{ij}_{l}$ exponentially. As a result, for $t>t_{0}$ we have $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<C\exp(-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(}t{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-t_{0}}))$, for some constant $C$. Case 2) $q^{*}_{k}=a^{ij}$, i.e., $q^{*}_{k}$ is anticoordinator-driven. Then, a time moment similar to $t_{0}$ defined in Case 1 exists such that for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, $|x_{l}(t)-a^{ij}_{l}|<C\exp(-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(}t{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-t_{0}}))$, for some constant $C$ and $t\geq t_{0}$. Therefore, in view of (18), it is straightforward to show that for any arbitrary small $\epsilon_{1}$, there exists some finite time $t_{1}+t_{0}$, such that $|x_{i}(t)-{q}^{*}_{k_{i}}|<\epsilon+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1}$ for $t>t_{1}+t_{0}$ resulting in $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<(\epsilon+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1}^{2}$. Having this and (i) we conclude that in both cases 1 and 2, for any arbitrary $\varepsilon$-neighborhood of $\mathbf{q}^{*}_{k}$, time $T>0$ can be found such that for all $t\geq T$, $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<\varepsilon$. As a result, the basin of attraction of point $\bm{q}^{*}_{k}$ equals $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})\}$. The basins of attraction of $\mathbf{q}^{*}_{1}$ and $\mathbf{q}^{*}_{\mathtt{Q}}$ can be obtained similarly. Part 2) When $x(0)=q^{*}_{k-1,k}$, where $q^{*}_{k-1,k}=o^{ij}$, two cases may happen. Case 1) The abstract state $x$ remains at $q^{*}_{k-1,k}$. Then for $l\neq j$, $\mathcal{V}_{l}(\bm{x})$ is a singleton and $\dot{x}_{l}$ is $\rho_{l}-x_{l}$ for $l\leq i$, or for $l>\mathsf{p}$ and $\mathsf{p}+\mathsf{p}^{\prime}+1-j<l$, and $\dot{x}_{l}$ is $-x_{l}$ for the remaining $l$’s. As a result, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{j\}$, the final value of $x_{l}$ is $o^{ij}_{l}$ and consequently $x_{l}$ will approach $o^{ij}_{l}$ exponentially. Then for any arbitrary small $\varepsilon$, in a finite time $T$ we have $|x_{l}(t)-o^{ij}_{l}|<\varepsilon$ for $t\geq T$ and subsequently $|x_{j}(t)-o^{ij}_{j}|<|x(t)-o^{ij}|+\Sigma_{l\neq j}|(x_{l}(t)-o^{ij}_{l})|\leq(\mathsf{p}+\mathsf{p}^{\prime}-1)\varepsilon$ resulting in $|\bm{x}(t)-\mathbf{o}^{ij}|\leq\varepsilon\sqrt{(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}-1)}$ for $t\geq T$. Hence, $\mathbf{o}^{ij}$ is the limit point. Case 2) The abstract state $x$ leaves $q^{*}_{k-1,k}$. Then, the provided reasoning for the case $x\neq q^{*}_{k-1,k}$ will be applicable resulting in $\mathbf{q}^{*}_{k-1}$ or $\mathbf{q}^{*}_{k}$ to be the limit point. Overall, the limit set of $\mathbf{q}^{*}_{k-1,k}$ is $\{\mathbf{q}^{*}_{k-1},\mathbf{q}^{*}_{k-1,k},\mathbf{q}^{*}_{k}\}$. ∎
VII-H Proof of Proposition 3

### v3

@@ VII-H Proof of Theorem 2
As for Part 1, the instability of coordinator-driven equilibria is shown by providing an initial condition from which for a small $\epsilon$, there exists no $\delta$ to satisfy the definition of stability. The stability of clean-cut and anticoordinator-driven equilibria is first shown for the abstract dynamics using Theorem 1 in [51]. Then, the definition of asymptotic stability is used to show that of clean-cut and anticoordinator-driven equilibria for the continuous-time population dynamics. Part 2 of the theorem is shown by using the definition of the limit set.
@@
Part 1) Regarding the instability analysis of a coordinator-driven equilibrium point, say $\mathbf{q}^{*}_{k-1,k}=\mathbf{o}^{ij}$, consider the following initial condition $\bm{x}(0)=\mathbf{o}^{ij}+\delta^{*}\mathbf{e}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}$ for an arbitrarily small $\delta^{*}\in(0,\min\{\tau^{\prime}_{j+1},\tau_{i}\}-\tau^{\prime}_{j})$. The corresponding abstract state $x(0)$ will then belong to $(\tau^{\prime}_{j},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$. Pick $\epsilon$ to be equal to $0.5(\min\{\tau^{\prime}_{j+1},\tau_{i},{\eta}_{i}+{\eta}_{j}^{\prime}\}-\tau^{\prime}_{j})$. In view of (7), at such $x(0)$, $\bm{\mathcal{V}}(\bm{x})$ is a singleton. Moreover, considering the structure of $\mathbf{o}^{ij}$, at $x(0)$ we have $\dot{x}_{l}=0$ for $l\in\{1,\ldots,\mathsf{p}+\mathsf{p}^{\prime}-j,\mathsf{p}+\mathsf{p}^{\prime}-j+2,\ldots,\mathsf{p}+\mathsf{p}^{\prime}\}$. Indeed, the equation $\dot{x}_{l}=0$ holds true as long as $x\in(\tau^{\prime}_{j},\min\{\tau^{\prime}_{j+1},\tau_{i}\})$. On the other hand, $x(0)>\tau^{\prime}_{j}$ and in view of (14), we have $\dot{x}_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)=\rho^{\prime}_{j}-x_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)>0$, for all $t\leq t^{\prime}$, where $t^{\prime}=\inf\{t\geq 0|x_{\mathsf{p}+\mathsf{p}^{\prime}-j+1}(t)=\min\{\alpha_{i,j},{\rho}^{\prime}_{j}\}\}$, where $\alpha_{i,j}=\min\{\tau^{\prime}_{j+1},\tau_{i}\}-({\eta}_{i}+{\eta}_{j-1})$. If $\min\{\alpha_{i,j},{\rho}^{\prime}_{j}\}=\alpha_{i,j}$, then $x(t^{\prime})=\min\{\tau^{\prime}_{j+1},\tau_{i}\}$. Otherwise, $x(t^{\prime})={\eta}_{j}+{\eta}_{i}$. In either case, we observe that $|x(t^{\prime})-\tau_{j}^{\prime}|$ is greater than the selected value for $\epsilon$, i.e., $0.5(\min\{\tau^{\prime}_{j+1},\tau_{i},{\eta}_{i}+{\eta}_{j}^{\prime}\}-\tau^{\prime}_{j})$, and the value of $\delta^{*}$ does not impact the value of $x(t^{\prime})$. This indicates instability of ${o}^{ij}$ and in turn $\mathbf{o}^{ij}$. The stability properties of the other two types of equilibrium points, are first shown for the abstract dynamics. Let $V_{k}(x):\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}=\frac{1}{2}(x-q^{*}_{k})^{2}$, where $q^{*}_{k}={\eta}_{i}+{\eta}^{\prime}_{j}$ if $q^{*}_{k}$ is clean-cut, i.e., $q^{*}_{k}=c^{ij}$ and $q^{*}_{k}=\tau_{i}$ if it is anticoordinator-driven, i.e., $q^{*}_{k}=a^{ij}$. The set-valued derivative of $V_{k}(x)$ w.r.t (11) will be $\mathcal{D}(x)=\{(x-q^{*}_{k})\nu|\nu\in\mathcal{X}(x)\}$ [51]. We also have $\bar{f}(x)=\max\{y|y\in\mathcal{X}(x)\}<0$ for $x\in(q^{*}_{k},q^{*}_{k,k+1})$ and $\underline{f}(x)=\min\{y|y\in\mathcal{X}(x)\}>0$ for $x\in(q^{*}_{k-1,k},q^{*}_{k})$ (refer to the proof of A2). Hence, $\max\mathcal{D}(x)<0$ for $x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})/\{q^{*}_{k}\}$. According to [51, Theorem 1], $q^{*}_{k}$ is a strongly asymptotically stable equilibrium point for (11). Also the basin of attraction of $q^{*}_{k}$, will be $(q^{*}_{k-1,k},q^{*}_{k,k+1})$. As for $q^{*}_{1}$ (resp. $q^{*}_{\mathtt{Q}}$), we have $\bar{f}(0)>0$ (resp. $\bar{f}(1)<0$) and this case can be handled similarly which results in $\mathcal{A}(q^{*}_{1})=[0,q^{*}_{1,2})$ (resp. $\mathcal{A}(q^{*}_{\mathtt{Q}})=(q^{*}_{\mathtt{Q}-1,\mathtt{Q}},1]$). Now, we show the asymptotic stability of $\mathbf{q}^{*}_{k}$ when it is an anticoordinator-driven equilibrium point, i.e., $\mathbf{q}^{*}_{k}=\mathbf{a}^{ij}$. The case where $\mathbf{q}^{*}_{k}$ is clean-cut can be handled similarly. We show that for every $\varepsilon>0$, there exists a $\delta>0$ such that if $|\bm{x}(0)-\mathbf{a}^{ij}|<\delta$, then $|\bm{x}(t)-\mathbf{a}^{ij}|<\varepsilon$ for all $t\geq 0$, and $|\bm{x}(t)-\mathbf{a}^{ij}|\rightarrow 0$ as $t\rightarrow\infty$. Consider a small enough $\epsilon>0$, such that at an $\epsilon$-neighborhood of $a^{ij}$, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, $\mathcal{V}_{l}(\bm{x})$ is a singleton and $\dot{x}_{l}=\rho_{l}-x_{l}$ for $l<i$ or for $l$ satisfying both $\mathsf{p}<l$ and $0<\mathsf{p}+\mathsf{p}^{\prime}+1-l\leq j$. For the remaining $l$’s, $\dot{x}_{l}=-x_{l}$. From the stability of the abstract dynamics, the existence of $\delta_{x}>0$ follows such that $|x(0)-a^{ij}|<\delta_{x}$ results in $|x(t)-a^{ij}|<\epsilon$ for $t\geq 0$. As a result, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, the final value of $x_{l}$ is $a^{ij}_{l}$ and consequently $x_{l}$ will approach $a^{ij}_{l}$ exponentially, i.e.,
@@
On the other hand, starting from $\delta_{x}$-neighborhood of $a^{ij}$, in view of (18), we have $|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<\big{(}(\mathsf{p}+\mathsf{p}^{\prime})\max_{l}|x_{l}(0)-a^{ij}_{l}|+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|e^{-t}\big{)}^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|^{2}e^{-2t}$ and, in turn, $|\bm{x}(t)-\mathbf{a}^{ij}|$ $<\max_{l}|x_{l}(0)-a^{ij}_{l}|\big{(}\mathsf{p}+\mathsf{p}^{\prime}+(\mathsf{p}+\mathsf{p}^{\prime}-1)e^{-t})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)e^{-2t}$$\big{)}^{0.5}.$ If $\delta$ is less than $\delta_{x}/(\mathsf{p}+\mathsf{p}^{\prime})$, then in view of the relation between Euclidean norm and Norm-1, $\sum_{l=1}^{\mathsf{p}+\mathsf{p}^{\prime}}|x_{l}(0)-a_{l}^{ij}|$ will be less than $\delta_{x}/\sqrt{\mathsf{p}+\mathsf{p}^{\prime}}$. and, in turn, $|x(0)-a^{ij}|<\delta_{x}$. Hence, $\mathbf{a}^{ij}$ is stable because by choosing $\delta=\min\{\frac{\delta_{x}}{\mathsf{p}+\mathsf{p}^{\prime}},\frac{\varepsilon}{\sqrt{(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}+(\mathsf{p}+\mathsf{p}^{\prime}-1))^{2}+(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}-1)}\}},$ we have $|x(t)-a^{ij}|<\epsilon$ for all $t\geq 0$, and consequently $|\bm{x}(t)-\mathbf{a}^{ij}|<\varepsilon$. As for asymptotic stability, we have $|\bm{x}(t)-\mathbf{a}^{ij}|^{2}<(|x(t)-a^{ij}|+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|e^{-t})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\max_{l}|x_{l}(0)-a^{ij}_{l}|^{2}e^{-2t}$, for $t\geq 0$. Since each term in the right hand side approaches zero, we have $\lim|\bm{x}(t)-\mathbf{a}^{ij}|\rightarrow 0$ as $t\rightarrow\infty$. Now, we obtain the basin of attraction of $\mathbf{q}^{*}_{k}$. (i) It is straightforward to show that if $x(0)$ falls in the basin of attraction $q^{*}_{k}$, the abstract dynamics will converge to an arbitrary small neighborhood of $q^{*}_{k}$ in a finite time. Now, depending on the types of $q^{*}_{k}$, two cases may happen: Case 1) $q^{*}_{k}=c^{ij}$, i.e., $q^{*}_{k}$ is clean-cut, then let $t=t_{0}$ be the time instant at which the abstract state $x$ enters a small enough $\epsilon$-neighborhood of $q^{*}_{k}$, such that for $l\leq i$ or $l\geq\max\{\mathsf{p}+\mathsf{p}^{\prime}-j+1,\mathsf{p}+1\}$, we have $\dot{x}_{l}=\rho_{l}-x_{l}$ and for the remaining $l$’s, we have $\dot{x}_{l}=-x_{l}$. Then for $t>t_{0}$, the asymptotic value of $x_{l}$ for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]$ is the same as the $l$-th entry of $\mathbf{q}^{*}_{k}$. Moreover, each $x_{l}$ approaches $c^{ij}_{l}$ exponentially. As a result, for $t>t_{0}$ we have $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<C\exp(-{(}t{-t_{0}}))$, for some constant $C$. Case 2) $q^{*}_{k}=a^{ij}$, i.e., $q^{*}_{k}$ is anticoordinator-driven. Then, a time moment similar to $t_{0}$ defined in Case 1 exists such that for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{i\}$, $|x_{l}(t)-a^{ij}_{l}|<C\exp(-{(}t{-t_{0}}))$, for some constant $C$ and $t\geq t_{0}$. Therefore, in view of (19), it is straightforward to show that for any arbitrary small $\epsilon_{1}$, there exists some finite time $t_{1}+t_{0}$, such that $|x_{i}(t)-{q}^{*}_{k_{i}}|<\epsilon+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1}$ for $t>t_{1}+t_{0}$ resulting in $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<(\epsilon+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1})^{2}+(\mathsf{p}+\mathsf{p}^{\prime}-1)\epsilon_{1}^{2}$. Having this and (i) we conclude that in both cases 1 and 2, for any arbitrary $\varepsilon$-neighborhood of $\mathbf{q}^{*}_{k}$, time $T>0$ can be found such that for all $t\geq T$, $|\bm{x}(t)-\mathbf{q}^{*}_{k}|<\varepsilon$. As a result, the basin of attraction of point $\bm{q}^{*}_{k}$ equals $\{\bm{x}\in\bm{\mathcal{X}}_{ss}|x\in(q^{*}_{k-1,k},q^{*}_{k,k+1})\}$. The basins of attraction of $\mathbf{q}^{*}_{1}$ and $\mathbf{q}^{*}_{\mathtt{Q}}$ can be obtained similarly. Part 2) When $x(0)=q^{*}_{k-1,k}$, where $q^{*}_{k-1,k}=o^{ij}$, two cases may happen. Case 1) The abstract state $x$ remains at $q^{*}_{k-1,k}$. Then for $l\neq j$, $\mathcal{V}_{l}(\bm{x})$ is a singleton and $\dot{x}_{l}$ is $\rho_{l}-x_{l}$ for $l\leq i$, or for $l>\mathsf{p}$ and $\mathsf{p}+\mathsf{p}^{\prime}+1-j<l$, and $\dot{x}_{l}$ is $-x_{l}$ for the remaining $l$’s. As a result, for $l\in[\mathsf{p}+\mathsf{p}^{\prime}]-\{j\}$, the final value of $x_{l}$ is $o^{ij}_{l}$ and consequently $x_{l}$ will approach $o^{ij}_{l}$ exponentially. Then for any arbitrary small $\varepsilon$, in a finite time $T$ we have $|x_{l}(t)-o^{ij}_{l}|<\varepsilon$ for $t\geq T$ and subsequently $|x_{j}(t)-o^{ij}_{j}|<|x(t)-o^{ij}|+\Sigma_{l\neq j}|(x_{l}(t)-o^{ij}_{l})|\leq(\mathsf{p}+\mathsf{p}^{\prime}-1)\varepsilon$ resulting in $|\bm{x}(t)-\mathbf{o}^{ij}|\leq\varepsilon\sqrt{(\mathsf{p}+\mathsf{p}^{\prime})(\mathsf{p}+\mathsf{p}^{\prime}-1)}$ for $t\geq T$. Hence, $\mathbf{o}^{ij}$ is the limit point. Case 2) The abstract state $x$ leaves $q^{*}_{k-1,k}$. Then, the provided reasoning for the case $x\neq q^{*}_{k-1,k}$ will be applicable resulting in $\mathbf{q}^{*}_{k-1}$ or $\mathbf{q}^{*}_{k}$ to be the limit point. Overall, the limit set of $\mathbf{q}^{*}_{k-1,k}$ is $\{\mathbf{q}^{*}_{k-1},\mathbf{q}^{*}_{k-1,k},\mathbf{q}^{*}_{k}\}$. ∎
VII-I Proof of Proposition 3